[Plate 3] FIGURE IV-6: The first 9 frames of Fig. IV-4, as viewed from above the orbital plane of the two galaxies (X_{o}, Y_{o}). The particles from the NW disk are color coded according to their position at T=8. Magenta represents the innermost, most tightly bound disk material and accounts for 74% of the original NW disk. Blue is the material next farthest out and accounts for 12% of the disk material, followed by green (8%) and finally red (6%). Initial, luminous tails composed of 27% of the disk material are raised (blue, green, and red, T=16), but most of this falls back in quickly (blue, by T=32) . At the present time, only the red particles remain in the tails, and these continue to rain back upon the remnant. (See page 140)

[Plate 4] FIGURE IV-8: (a) The radial period versus pericentric distance for the tail particles, calculated according to their energies and angular momenta at T=72 (0.6 h^{−1} Gyr). The particles are color coded according to their return times and radii of closest approach (R_{peri}). The large black dot is drawn at the location of the observed clump in the NW tail (HGvGS), which is predicted to fall to within 14 h^{−1} kpc in 3.1 h^{−1} Gyr. (b) The sky view of the simulation at T=72, with the tail particles color coded as in (a). This figure is used to construct a “mask” used on the observational data to derived the quantities listed in Table IV-1. (See page 144)

[Plate 5] FIGURE AII-6: Calibrated B-R color map of the Arp 295 system. (See page 198)

[Plate 6] FIGURE AIV-7: Calibrated B-R color map of the NGC 520 system. (See page 214)

[Plate 7] FIGURE AVI-2: Calibrated B-R color map of the NGC 7252 system. (See page 233)

[Plate 8] FIGURE AVI-3: Uncalibrated B-R color map of the NGC 4038/9 system. (See page 233)

[Plate 9] FIGURE AVIII-4: NW disk at periapse, looking normal to the surface, and using the same color coding as in Fig. IV-6. The x-axis is aligned along the line of approach of the two disks. Rotation is counter-clockwise in this view. Material that winds up at different distances along the tails is drawn from slices of the progenitor disk at orbital periapse. (See page 249)

[Plate 10] FIGURE VIII-5: E − V_{x} plane of the NW disk at periapse, using the same color coding as in Fig. AVIII-4. Tails are drawn from the highest energy material which is moving away from the companion. (See page 249)